# From Feedback to Checklists: Grounded Evaluation of AI-Generated Clinical Notes

### v1

@@ F.2 Evaluation
For simplicity, we set up the task as a binary classification task, by using data with star ratings of 1 as negative and 5 as positive (1113 and 680 samples respectively). We can also set up the task as a multiclass classification task, by using the star ratings as the classes. We then train a Logistic Regression model with 5-fold cross-validation on 80% of the feedback and test with the remaining 20%.
@@
From Fig. 7, we see that the baseline checklist has an average unit test scores of 0.92 over 10 questions, while the feedback checklist has an average unit test score of 0.92 over 25 questions.
@@
Appendix G Challenges & Limitations
As described in §4, we made several simplifying design choices in our pipeline to generate a checklist from user feedback. Checklists for other note sections have not been implemented yet. The choice to exclude “N/A” as an answer option simplifies evaluation of checklists, but restricts the space of useful questions. Likewise, dropping questions that compare across note sections may filter out useful questions. Our process of deduplicating questions has not been rigorously evaluated, and we may be losing nuanced questions from this step. We have not yet implemented a dynamic pipeline, where the checklist is updated as new feedback comes in.
Our premise of grounding checklist questions in user feedback, means that aspects of notes that do not appear in feedback may not be covered in our questions. E.g., if there are no complaints about pronoun usage, there may not be a checklist question about using correct pronouns. Our checklist items are also yet not ordered or weighted in scoring, which is a feature that could be useful (Wei et al., 2025).
We rely heavily on LLMs in our pipeline and evaluation, which introduces its own set of challenges. When using LLMs-as-a-Judge, there is a lack of ground truth labels to determine whether clinical notes satisfy checklist questions, such as for ACI-bench notes. For unit tests, we have not verified that the LLMs correctly rewrite reference notes to fail the checklist questions. Many refinement steps involve the use of LLMs for tagging, but these have not been thoroughly verified. Each checklist question essentially represents its own distinct task.
Evaluation is also a difficult problem for this task, since each checklist question can be a substantial problem on its own. As such, our current metrics are not exhaustive. Reference notes from ACI-Bench vary significantly in quality, meaning they may not all achieve perfect checklist scores. Finally, we rely on proxies for positive signals, such as internal pairwise preference data, though these proxies also have their own limitations (e.g., poor agreement between raters).
Appendix H Future Work
@@
We can further try tuning LLM-as-a-Judge specifically for our task of checklist scoring. While the current setup simply has the LLM-evaluator produce a “Yes”/“No” answer for each checklist question, we can also allow the LLM to reason about the note and checklist question and provide a more detailed explanation of its score.

### v2

@@ F.2 Evaluation
For simplicity, we set up the task as a binary classification task, by using data with star ratings of 1 as negative and 5 as positive (1113 and 680 samples respectively). We then train a Logistic Regression model with 5-fold cross-validation on 80% of the feedback and test with the remaining 20%.
@@
From Fig. 7, we see that the baseline checklist has an average unit test scores of 0.82 over 10 questions, while the feedback checklist has an average unit test score of 0.89 over 20 questions.
@@
Appendix G Future Work
@@
We can further try tuning LLM-as-a-Judge specifically for our task of checklist scoring. While the current setup simply has the LLM-evaluator produce a “Yes”/“No” answer for each checklist question, we can also allow the LLM to reason about the note and checklist question and provide a more detailed explanation of its score. Allowing the evaluator to cross-validate the note with structured EHR data could enhance accuracy by catching factual inconsistencies that text-only evaluation misses.